# Accelerated Convergence of Frank–Wolfe Algorithms with Adaptive Bregman Step-Size Strategy

### v1

@@ 2.1 Bregman Distances
Definition 2.4 (Triangle Scaling Property [Hanzely2021-sz, Definition 2]).
@@
$D_{\phi}((1-\gamma)x+\gamma y,(1-\gamma)x+\gamma z)\leq\gamma^{\nu}D_{\phi}(y,z).$ (2.2)
Now, substituting $z\leftarrow x$ on the left-hand side of (2.2), we obtain
@@
We will now show that a stronger version can be obtained if $\phi$ is strictly convex, so that we can rephrase $\nu$ with $1+\nu$ where $\nu\in(0,1]$, i.e., the right-hand side is superlinear, which will be crucial for the convergence analysis later.
Lemma 2.5.
Given a kernel generating distance $\phi\in\mathcal{G}(C)$, if $\phi$ is strictly convex, then there exists $\nu\in(0,1]$ such that, for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, it holds that
@@
Obviously, (2.3) holds if $y=x$ or $\gamma\in\{0,1\}$. In what follows, we assume $y\neq x$ and $10<\gamma<$ 1. Let $g(\nu):=\gamma^{1+\nu}D_{\phi}(y,x)-D_{\phi}((1-\gamma)x+\gamma y,x)$ for all $x,y\in\operatorname{int}\operatorname{dom}\phi$, $y\neq x$ and all $\gamma\in(0,1)$. Using $D_{\phi}(y,x)>0$ due to the strict convexity of $\phi$, we have, for any $\nu\geq 0$,
@@
In the remainder of the paper, if not stated otherwise, we will assume that $\phi$ is strictly convex and as such for a given $\phi$ and $C$ there exists a $\nu\in(0,1]$ such that $D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x)$ holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$.
2.2 Smooth Adaptable Property
Consider a pair of functions $(f,\phi)$ satisfying the following conditions:
  1. (i)
    $\phi\in\mathcal{G}(C)$,
  2. (ii)
    $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C\equiv\operatorname{int}\operatorname{dom}\phi$.
The pair $(f,\phi)$ is said to be $L$-smooth adaptable (for short: $L$-smad) on $C$ if there exists $L>0$ such that $L\phi-f$ and $L\phi+f$ are convex on $C$.
@@ 2.2 Smooth Adaptable Property
Lemma 2.7 (Extended Descent Lemma [Bolte2018-zt, Lemma 2.1]).
@@
From this, it can be seen that the $L$-smad property for $(f,\phi)$ provides upper and lower approximations for $f$ majorized by $\phi$ with $L>0$. In addition, if $\phi(x)=\frac{1}{2}\|x\|^{2}$ on $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$, Lemma 2.7 corresponds to the classical descent lemma. While the $L$-smad property might seem unfamiliar at first, it is a natural generalization of $L$-smoothness, and examples of functions $f$ and $\phi$ satisfying the $L$-smad property are given, e.g., in [Bauschke2017-hg, Lemmas 7 and 8][Bolte2018-zt, Lemma 5.1], [Dragomir2021-rv, Propositions 2.1 and 2.3], [Mukkamala2019-mk, Proposition 2.1], [Takahashi2024-wz, Proposition 24], [Takahashi2023-uh, Theorem 1], and [Takahashi2024-cu, Theorem 15].
@@
Lemma 2.8 (Primal progress from the smooth adaptable property).
Let the pair of functions $(f,\phi)$ be $L$-smad with a strictly convex function $\phi$ and let $x_{+}=(1-\gamma)x+\gamma v$ with $x,v\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$. Then it holds:
@@
Using Lemma 2.7 and substituting $x_{+}$ for $x$ and $x$ for $y$, we have
@@
It holds that $D_{\phi}(x_{+},x)\leq\gamma^{1+\nu}D_{\phi}(v,x)$ for some $\nu\in(0,1]$ by Lemma 2.5. Therefore, this provides the desired inequality. ∎
Often the point $v\in P\subset\operatorname{int}\operatorname{dom}\phi$ in Lemma 2.8 is chosen as a Frank–Wolfe vertex, i.e.,
$v\in\operatorname*{argmax}_{u\in P}\langle\nabla f(x),x-u\rangle,$
but other choices, e.g., those arising from away-directions from the away-step FW method, are also possible, as we will see in Section 3.2.
2.3 Weakly Convex Functions
A function $f$ is called weakly convex if it becomes convex upon adding a quadratic perturbation.
Definition 2.9 (Weakly Convex Function [Davis2018-bb, Nurminskii1975-vp]).
A function $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is said to be $\rho$-weakly convex for some $\rho>0$ if the function $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex.
@@ 2.3 Weakly Convex Functions
For examples of such $f$, see [Davis2018-bb, Examples 3.1 and 3.2]. Moreover, it is easy to see that any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex [Vial1983-oo, Proposition 4.11] and [Wu2007-dt], the rationale being that we can simply shift the function by the quadratic belonging to its smallest eigenvalue. Since the proof of this fact is omitted therein, we provide the argument for the sake of completeness below.
Proposition 2.10.
@@
Let $g=\lambda_{\min}(\nabla^{2}f(\cdot))$. Obviously, $g$ is continuous because $\lambda_{\min}(\cdot)$ and $\nabla^{2}f(\cdot)$ are continuous. Therefore, there exists the minimum value of $g$ on $P$ due to the continuity of $g$ and the compactness of $P$. For $\rho=|\min_{x\in P}g(x)|$, we have $\nabla^{2}f(x)+\rho I\succeq O$ for any $x\in P$, which implies $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex on $P$, i.e., $f$ is $\rho$-weakly convex on $P$. ∎
Minimizing a nonconvex $\mathcal{C}^{2}$-function over a compact set $P$ is equivalent to minimizing a weakly convex function over $P$. However, it is important to note that when $\rho$ becomes large, the assumptions of algorithms may not be satisfied (see also Theorems 5.5 and 5.9).
2.4 Growth Conditions
In [garber2015faster] (see also [Garber2020-tr]) the linear convergence of the FW algorithm for convex optimization under the quadratic growth condition, which is a weaker assumption than assuming strong convexity was established over strongly convex sets and later generalized to uniformly convex sets in [KAP2020] as well as to conditions weaker than quadratic growth in [KDP2018, KDP2018jour]. Local variants of these notions, as necessary, e.g., for the nonconvex case, have been studied in [KAP2021] in the context of Frank–Wolfe methods.
Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be a proper lower semicontinuous function and $P\subset\mathbb{R}^{n}$ be a compact convex set. Let $\mathcal{X}^{*}\neq\emptyset$ be the set of optimal solutions, i.e., $\mathcal{X}^{*}:=\operatorname*{argmin}_{x\in P}f(x)$, and let $f^{*}=\min_{x\in P}f(x)$ and $\zeta>0$. The function $f$ satisfies the Hölder error bound condition on $P$ if there exists constants $q\geq 1$ and $\mu>0$ such that, for any $x\in P$,
The Hölder error bound condition shows sharpness bounds on the primal gap (see, e.g., [bolte17, alex17sharp]), which has also been extensively analyzed for Frank–Wolfe methods in [KDP2018, KDP2018jour]. The convergence analysis for nonconvex optimization under the quadratic growth or Hölder error bound condition was established in [Davis2024-rm, Davis2018-bb, Davis2024-og]. We distinguish between the global condition and the local condition of Definition 2.11. We employ the Hölder error bound condition for convex optimization while we use the local quadratic growth for nonconvex optimization.
Next, we show in the following lemma that the Hölder error bound condition immediately provides a bound on the primal optimality gap, which will be useful for establishing convergence rates of the proposed algorithms as well as the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks], provided $f$ is convex.
Lemma 2.12 (Primal gap bound from Hölder error bound).
Let $f$ be a convex function and satisfy the Hölder error bound condition with $q\geq 1$ and $\mu>0$. Let $x^{*}$ be the unique minimizer of $f$ over $P$. Then, the following argument holds in general, for all $x\in P$:
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}},$ (2.5)
@@ 2.4 Growth Conditions
By first applying convexity and then the Hölder error bound condition for any $x^{*}\in\mathcal{X}^{*}$ with $f^{*}=f(x^{*})$ it holds:
@@
Remark 2.13.
With the remark from above, we can immediately relate the Hölder error bound condition to the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks]; the Łojasiewicz inequality is used to establish convergence analysis [Bolte2007-xr]. To this end, using (2.5), we estimate $\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\leq\frac{\|\nabla f(x)\|\|x-x^{*}\|}{\|x-x^{*}\|}=\|\nabla f(x)\|$ to then obtain the weaker condition:
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\|\nabla f(x)\|^{\frac{q}{q-1}}=c^{\frac{1}{\theta}}\|\nabla f(x)\|^{\frac{1}{\theta}},$ (2.6)
where $c=\left(\frac{q}{\mu}\right)^{\frac{1}{q}}$ and $\theta=\frac{q-1}{q}$. Inequality (2.6) is called the $c$-Łojasiewicz inequality with $\theta\in[0,1)$. If $\mathcal{X}^{*}\subseteq\operatorname{ri}P$, then the two conditions are equivalent. However, if the optimal solution(s) are on the boundary of $P$ as is not infrequently the case, then the two conditions are not equivalent as $\|\nabla f(x)\|$ might not vanish for $x\in\mathcal{X}^{*}$, whereas $\langle\nabla f(x),x-x^{*}\rangle$ does, i.e., the Hölder error bound condition is tighter than the one induced by the Łojasiewicz inequality.
The next lemma shows that we can also obtain primal gap bounds in the weakly convex case together with (local) quadratic growth. Its proof can be found in Appendix A.1.
Lemma 2.14 (Primal gap bound from the quadratic growth).
@@
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$ (2.7)
@@
Remark 2.15.
In the same vein as the discussion in Remark 2.13, we can immediately relate the local $\mu$-quadratic growth condition with $\zeta>0$ to the local Polyak–Łojasiewicz (PL) inequality. Using (2.7), for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\|\nabla f(x)\|^{2},$ (2.8)
which is equivalent to the PL inequality [Polyak1963-dr], also called the gradient dominance property [Combettes2020-jj], with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$. The PL inequality is also equivalent to the Łojasiewicz inequality with $\theta=\frac{1}{2}$. Strongly convex functions satisfy the PL inequality [Braun2022-ew, Lemma 2.13]. Note that we will not have the primal gap bound under the Hölder error bound condition and weak convexity because an inequality like (A.1) does not hold from the Hölder error bound condition.
2.5 Scaling Inequalities, Geometric Hölder Error Bounds, and Contractions
@@
Proposition 2.16 (Scaling inequality for inner optima from convexity [Braun2022-ew, Proposition 2.16]).
@@
Proposition 2.17 (Scaling inequality for inner optima from weak convexity).
@@
$\langle\nabla f(x),v\rangle\leq\langle\nabla f(x),x^{*}-rz\rangle=\langle\nabla f(x),x^{*}\rangle-r\|\nabla f(x)\|.$ (2.9)
@@
By rearranging (2.9) and using $\langle\nabla f(x),x-x^{*}\rangle\geq 0$, we obtain
@@
The authors in [Lacoste-Julien2015-gb] defined a geometric distance-like constant of a polytope, known as the pyramidal width, to analyze the convergence of the away-step Frank–Wolfe algorithm (and other variants that use the pyramidal width) over polytopes. It can be interpreted as the minimal $\delta>0$ satisfying the following scaling inequality (2.10), which plays a central role in establishing convergence rates for the away-step FW algorithm; see [Braun2022-ew] for an in-depth discussion.
Lemma 2.18 (Scaling inequality via pyramidal width [Braun2022-ew, Theorem 2.26][Lacoste-Julien2015-gb, Theorem 3]).
Let $P\subset\mathbb{R}^{n}$ be a polytope and let $\delta$ denote the pyramidal width of $P$. Let $x\in P$, and let $\mathcal{S}$ denote any set of vertices of $P$ with $x\in\operatorname{conv}\mathcal{S}$. Let $\psi$ be any vector, so that we define $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\psi,v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\psi,v\rangle$. Then for any $y\in P$
$\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle\geq\delta\frac{\langle\psi,x-y\rangle}{\|x-y\|}.$ (2.10)
Note that Lemma 2.18 does not require the convexity of $f$, and we will use it for nonconvex optimization. When $\phi$ is $\sigma$-strongly convex and $\langle\psi,x-y\rangle\geq 0$, Lemma 2.18 implies
@@
We will now introduce the more compact notion of geometric Hölder error bound condition, which simply combines the pyramidal width and the Hölder error bound condition of the function $f$; see also [Braun2022-ew, Lemma 2.27] for details when $q=2$.
Lemma 2.19 (Geometric Hölder Error Bound).
Let $P$ be a polytope with pyramidal width $\delta>0$ and let $f$ be a convex function and satisfy the Hölder error bound condition with $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\nabla f(x),v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x),v\rangle$ with $\mathcal{S}\subseteq\operatorname{Vert}{P}$, so that $x\in\operatorname{conv}{\mathcal{S}}$, we have
@@
Combining (2.10) with Lemma 2.12 for $\psi=\nabla f(x)$, we have
@@
Lemma 2.20 (From contractions to convergence rates [Braun2022-ew, Lemma 2.21]).

### v3

@@ A.1 Properties of Bregman Distances
Definition A.2 (Triangle Scaling Property [Hanzely2021-sz, Definition 2]).
@@
$D_{\phi}((1-\gamma)x+\gamma y,(1-\gamma)x+\gamma z)\leq\gamma^{\nu}D_{\phi}(y,z).$ (A.1)
Now, substituting $z\leftarrow x$ on the left-hand side of (A.1), we obtain
@@
We will now show that a weaker version of the triangle scaling property can be obtained if $\phi$ is strictly convex, so that we can rephrase $\nu$ with $1+\nu$ where $\nu>0$, i.e., the right-hand side is superlinear.
Lemma A.3.
Given a kernel generating distance $\phi\in\mathcal{G}(C)$, if $\phi$ is strictly convex, then for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, there exists $\nu>0$ such that, it holds that
$D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x).$ (A.2)
@@
Obviously, (A.2) holds if $y=x$ or $\gamma\in\{0,1\}$. In what follows, we assume $y\neq x$ and $10<\gamma<$ 1. Let $g(\nu):=\gamma^{1+\nu}D_{\phi}(y,x)-D_{\phi}((1-\gamma)x+\gamma y,x)$ for all $x,y\in\operatorname{int}\operatorname{dom}\phi$, $y\neq x$ and all $\gamma\in(0,1)$. Using $D_{\phi}(y,x)>0$ due to the strict convexity of $\phi$, we have, for any $\nu\geq 0$,
@@
A.2 Lemmas of Smooth Adaptable Property
@@
Lemma A.4 (Extended Descent Lemma [Bolte2018-zt, Lemma 2.1]).
@@
From this, it can be seen that the $L$-smad property for $(f,\phi)$ provides upper and lower approximations for $f$ majorized by $\phi$ with $L>0$. In addition, if $\phi(x)=\frac{1}{2}\|x\|^{2}$ on $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$, Lemma A.4 corresponds to the classical descent lemma. While the $L$-smad property might seem unfamiliar at first, it is a natural generalization of $L$-smoothness, and examples of functions $f$ and $\phi$ satisfying the $L$-smad property are given, e.g., in [Bauschke2017-hg, Lemmas 7 and 8][Bolte2018-zt, Lemma 5.1], [Dragomir2021-rv, Propositions 2.1 and 2.3], [Mukkamala2019-mk, Proposition 2.1], [Takahashi2024-wz, Proposition 24], [Takahashi2023-uh, Theorem 1], and [Takahashi2024-cu, Theorem 15].
@@
Lemma A.5 (Primal progress from the smooth adaptable property).
Let the pair of functions $(f,\phi)$ be $L$-smad with a strictly convex function $\phi\in\mathcal{G}(C)$ and let $x_{+}=(1-\gamma)x+\gamma v$ with $x,v\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$. Then it holds:
@@
Using Lemma A.4 and substituting $x_{+}$ for $x$ and $x$ for $y$, we have
@@
It holds that $D_{\phi}(x_{+},x)\leq\gamma^{1+\nu}D_{\phi}(v,x)$ for some $\nu>0$ by Lemma A.3. Therefore, this provides the desired inequality. ∎
We also have a progress lemma from the Bregman short step-size.
Lemma A.6 (Progress lemma from the Bregman short step-size).
Proof.
Substituting $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$ with $x_{t},v_{t}\in P$ for $x_{+}$ in Lemma 2.4, we have
$=\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle,$
Appendix B Properties from Sharpness
B.1 Results from Weak Convexity
@@
Any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex [Vial1983-oo, Proposition 4.11] and [Wu2007-dt]. Since the proof of this fact is omitted therein, we provide the argument for the sake of completeness below.
Proposition B.1.
@@
Let $g=\lambda_{\min}(\nabla^{2}f(\cdot))$, where $\lambda_{\min}(\cdot)$ is the smallest eigenvalue. Obviously, $g$ is continuous because $\lambda_{\min}(\cdot)$ and $\nabla^{2}f(\cdot)$ are continuous. Therefore, there exists the minimum value of $g$ on $P$ due to the continuity of $g$ and the compactness of $P$. For $\rho=|\min_{x\in P}g(x)|$, we have $\nabla^{2}f(x)+\rho I\succeq O$ for any $x\in P$, which implies $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex on $P$, i.e., $f$ is $\rho$-weakly convex on $P$. ∎
Minimizing a nonconvex $\mathcal{C}^{2}$-function over a compact set $P$ is equivalent to minimizing a weakly convex function over $P$. However, it is important to note that when $\rho$ becomes large, the assumptions of algorithms may not be satisfied (see also Theorems D.3 and 5.3).
We have the key lemma for convergence analysis from weak convexity.
Lemma B.2 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex functions).
Let $x^{*}\in P$ an optimal solution of (1.1) and let $f^{*}=f(x^{*})$. Suppose that Assumptions 3.1iiiv, and that $f$ is $\rho$-weakly convex. Then for all $x\in P$, it holds:
$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (B.1)
Proof.
B.2 Primal Gap Bound
We show in the following lemma that the HEB condition immediately provides a bound on the primal optimality gap, which will be useful for establishing convergence rates of the proposed algorithms as well as the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks], provided $f$ is convex.
Lemma B.3 (Primal gap bound from Hölder error bound).
Let $f$ be a convex function and satisfy the HEB condition with $q\geq 1$ and $\mu>0$. Let $x^{*}$ be the unique minimizer of $f$ over $P$. Then, the following argument holds in general, for all $x\in P$:
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{\frac{q}{q-1}},$ (B.2)
@@ B.2 Primal Gap Bound
By first applying convexity and then the HEB condition for any $x^{*}\in\mathcal{X}^{*}$ with $f^{*}=f(x^{*})$ it holds:
@@
Remark B.4.
With the remark from above, we can immediately relate the HEB condition to the Łojasiewicz inequality [Bolte2007-xr, Lojasiewicz1963-sm, Lojasiewicz1965-ks]; the Łojasiewicz inequality is used to establish convergence analysis [Bolte2007-xr]. To this end, using (B.2), we estimate $\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\leq\frac{\|\nabla f(x)\|\|x-x^{*}\|}{\|x-x^{*}\|}=\|\nabla f(x)\|$ to then obtain the weaker condition:
$f(x)-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\|\nabla f(x)\|^{\frac{q}{q-1}}=c^{\frac{1}{\theta}}\|\nabla f(x)\|^{\frac{1}{\theta}},$ (B.3)
where $c=\left(\frac{q}{\mu}\right)^{\frac{1}{q}}$ and $\theta=\frac{q-1}{q}$. Inequality (B.3) is called the $c$-Łojasiewicz inequality with $\theta\in[0,1)$. If $\mathcal{X}^{*}\subseteq\operatorname{ri}P$, then the two conditions are equivalent. However, if the optimal solution(s) are on the boundary of $P$ as is not infrequently the case, then the two conditions are not equivalent as $\|\nabla f(x)\|$ might not vanish for $x\in\mathcal{X}^{*}$, whereas $\langle\nabla f(x),x-x^{*}\rangle$ does, i.e., the HEB condition is tighter than the one induced by the Łojasiewicz inequality.
The next lemma shows that we can also obtain primal gap bounds in the weakly convex case together with (local) quadratic growth.
Lemma B.5 (Primal gap bound from the quadratic growth).
@@
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\left(\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\right)^{2},$ (B.4)
@@
Proof.
$f(x)-f^{*}$ $=f(x)-f(x^{*})$
$\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}$
which implies
$\left(1-\frac{\rho}{\mu}\right)\left(f(x)-f^{*}\right)\leq\langle\nabla f(x),x-x^{*}\rangle=\frac{\langle\nabla f(x),x-x^{*}\rangle}{\|x-x^{*}\|}\|x-x^{*}\|.$ (B.5)
Using the local quadratic growth condition again, we have
which provides
or equivalently
Remark B.6.
In the same vein as the discussion in Remark B.4, we can immediately relate the local $\mu$-quadratic growth condition with $\zeta>0$ to the local Polyak–Łojasiewicz (PL) inequality. Using (B.4), for all $x\in[f\leq f^{*}+\zeta]\cap P$, we have
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}}\|\nabla f(x)\|^{2},$ (B.6)
which is equivalent to the PL inequality [Polyak1963-dr], also called the gradient dominance property [Combettes2020-jj], with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$. The PL inequality is also equivalent to the Łojasiewicz inequality with $\theta=\frac{1}{2}$. Strongly convex functions satisfy the PL inequality [Braun2022-ew, Lemma 2.13]. Note that we will not have the primal gap bound under the HEB condition and weak convexity because an inequality like (B.5) does not hold from the HEB condition.
B.3 Scaling Inequalities, Geometric Hölder Error Bounds, and Contractions
@@
Proposition B.7 (Scaling inequality from convexity when $x^{*}\in\operatorname{int}P$ [Braun2022-ew, Proposition 2.16]).
@@
Proposition B.8 (Scaling inequality from weak convexity when $x^{*}\in\operatorname{int}P$).
@@
$\langle\nabla f(x),v\rangle\leq\langle\nabla f(x),x^{*}-rz\rangle=\langle\nabla f(x),x^{*}\rangle-r\|\nabla f(x)\|.$ (B.7)
@@
By rearranging (B.7) and using $\langle\nabla f(x),x-x^{*}\rangle\geq 0$, we obtain
@@
The authors in [Lacoste-Julien2015-gb] defined a geometric distance-like constant of a polytope, known as the pyramidal width, to analyze the convergence of the away-step Frank–Wolfe algorithm (and other variants that use the pyramidal width) over polytopes. It can be interpreted as the minimal $\delta>0$ satisfying the following scaling inequality (B.8), which plays a central role in establishing convergence rates for the away-step FW algorithm; see [Braun2022-ew] for an in-depth discussion.
Lemma B.9 (Scaling inequality via pyramidal width [Braun2022-ew, Theorem 2.26][Lacoste-Julien2015-gb, Theorem 3]).
Let $P\subset\mathbb{R}^{n}$ be a polytope and let $\delta$ denote the pyramidal width of $P$. Let $x\in P$, and let $\mathcal{S}$ denote any set of vertices of $P$ with $x\in\operatorname{conv}\mathcal{S}$, where $\operatorname{conv}\mathcal{S}$ denotes the convex hull of $\mathcal{S}$. Let $\psi$ be any vector, so that we define $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\psi,v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\psi,v\rangle$. Then for any $y\in P$
$\langle\psi,v^{\textup{A}}-v^{\textup{FW}}\rangle\geq\delta\frac{\langle\psi,x-y\rangle}{\|x-y\|}.$ (B.8)
Note that Lemma B.9 does not require the convexity of $f$, and we will use it for nonconvex optimization. When $\phi$ is $\sigma$-strongly convex and $\langle\psi,x-y\rangle\geq 0$, Lemma B.9 implies
@@
We will now introduce the more compact notion of geometric Hölder error bound condition, which simply combines the pyramidal width and the HEB condition of the function $f$; see also [Braun2022-ew, Lemma 2.27] for details when $q=2$.
Lemma B.10 (Geometric Hölder Error Bound).
Let $P$ be a polytope with pyramidal width $\delta>0$ and let $f$ be a convex function and satisfy the HEB condition with $v^{\textup{FW}}=\operatorname*{argmin}_{v\in P}\langle\nabla f(x),v\rangle$ and $v^{\textup{A}}=\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x),v\rangle$ with $\mathcal{S}\subseteq\operatorname{Vert}{P}$, so that $x\in\operatorname{conv}{\mathcal{S}}$, we have
@@
Combining (B.8) with Lemma B.3 for $\psi=\nabla f(x)$, we have
@@
Lemma B.11 (From contractions to convergence rates [Braun2022-ew, Lemma 2.21]).